Micah wanted to buy a new iPod. He has $250 but must save 3% of it for his little brother's birthday present. What is the most expensive Ipod he can buy given the state tax on ipods is 7%?
Answer by mananth(16946)

Micah has $250
3% for gift
0.03 * 250 = $7.50
Balance amount = 250-7.50=242.5
...
Including 7% tax he can purchse an ipod worth $242.50
..
If the ipod costs $100 the sale price is 93
if the sale price has to be $242.50 then
..
93/100 * 242.50 should be the sale price
..
the pre-tax price should not be more than 225.50
..
m.ananth@hotmail.ca